# Computing hydration free energies of small molecules with first principles accuracy

### v2

@@ Abstract
Free energies play a central role in characterising the behaviour of chemical systems and are among the most important quantities that can be calculated by molecular dynamics simulations. The free energy of hydration in particular is a well-studied physicochemical property of drug-like molecules and is commonly used to assess and optimise the accuracy of nonbonded parameters in empirical forcefields, and also as a fast-to-compute surrogate of performance for protein-ligand binding free energy estimation. Machine learned potentials (MLPs) show great promise as more accurate alternatives to empirical forcefields, but are not readily decomposed into physically motivated functional forms, which has thus far rendered them incompatible with standard alchemical free energy methods that manipulate individual pairwise interaction terms. However, since the accuracy of free energy calculations is highly sensitive to the forcefield, this is a key area in which MLPs have the potential to address the shortcomings of empirical forcefields. In this work, we introduce an efficient alchemical free energy method compatible with MLPs, enabling, for the first time, calculations of biomolecular free energy differences with ab initio accuracy. Using a pretrained, transferrable, alchemically equipped MACE model, we demonstrate sub-chemical accuracy for the hydration free energies of organic molecules.
@@ I Introduction
Free energies are arguably the most important quantities accessible by molecular simulation in soft matter and play an important role in computational drug discovery [1, 2]. Free energy methods rely on rigorous thermodynamic protocols, enabling prediction of protein-ligand binding [3, 4], small molecule solvation and solubility [5], crystal polymorphs of organic molecules[6] and protein residue mutation [7].
@@
Whilst various methods for calculating free energies for small molecules via MD have been proposed, alchemical transformations have emerged as the de facto standard approach, allowing free energy estimates to be obtained in 1-2 GPU hours per compound [11, 12] Rather than explicitly simulating the a compound undergoing a change of state as it would happen in reality (which would be computationally prohibitive), alchemical calculations take advantage of the fact that the free energy is a state function and so the free energy difference between two state points is independent of the pathway connecting them. By introducing an alchemical parameter, $\lambda$, a new Hamiltonian is constructed as a linear combination of the Hamiltonians describing the two end states,
$H(\vec{r},\lambda)=\lambda H_{1}(\vec{r},\lambda)+(1-\lambda)H_{0}(\vec{r},\lambda)$ (1)
The free energy difference associated with the transformation can then be computed, for example by thermodynamic integration
@@
where $\langle\dots\rangle_{\lambda}$ denotes the ensemble average corresponding to the Hamiltonian $H(\lambda)$. Using such artificial pathways leads to enormous computational savings, since the sampling required to explore intermediate state points is orders of magnitude less than it would be for a real pathway describing the physical process. In practice, the transformation is accomplished by identifying the atoms that change during the transformation (i.e. those atoms whose chemical element changes, or that disappear entirely) and interpolating the corresponding forcefield parameters between the end states. This requires the use of so-called dummy atoms to model atoms that disappear during the transformation. These atoms maintain their bonded terms but have their interactions with the surroundings switched off. This allows for transformations like morphing a hydrogen atom to a methyl group or isolating a molecule from its environment, while minimally perturbing the phase space required to be explored by the system.
Early attempts to perform free energy calculations with MD faced energy conservation and convergence issues due to the divergence of the nonbonded energy (usually a Lennard-Jones potential) when partially decoupled atoms overlap [13, 14]. Beutler et al were the first to address this issue by incorporating softening parameters to scale nonbonded interactions as a function of the alchemical parameter $\lambda$, thus averting singularities in the potential energy as atoms come into contact [15]
@@
where $\epsilon$ and $\sigma$ are the Lennard-Jones well depth and radius, respectively, and $\alpha_{LJ}$ and $m$ are positive, tunable constants that control the smoothness of the softcore function. Although Beutler-type soft core potentials are commonly used in various simulation packages, it is worth noting that several alternative approaches have been suggested to further improve numerical stability and address specific failure modes [16, 17], and alternative functional forms with a more natural soft core have also been tried [18].
@@
While soft-core potentials have been widely successful, a significant limitation of current state-of-the-art free energy calculations lies in the accuracy of the empirical forcefields. Specifically, the constraints imposed by the functional form, omission of energetic contributions and restricted atom-type parameters in small molecule forcefields have a significant impact on accuracy [1]. Particularly notable are the widespread use of fixed-charge electrostatic models that do not account for geometry-dependent polarization, and the limited accuracy of torsional barriers, which often require refitting to bespoke DFT calculations of each new compound to achieve reasonable accuracy [19, 20, 21].
@@
In recent years, transferable machine learned potentials (MLP)s have been introduced as a compelling alternative to empirical forcefields, demonstrating significant advances in materials modelling and biomolecular simulation [22, 23, 24, 25, 26, 27]. Emerging from Behler and Parinello’s initial work [28], the field has expanded rapidly, with many architectures being proposed to accurately model the QM potential energy surface. A particularly successful innovation has been the specific encoding of permutational, rotatational and translational symmetries [29]. These approaches have led to a series of data-efficient, universal potentials for biomolecular and materials simulation, enabling accurate prediction of atomistic and thermodynamic properties for a wide range of chemical compositions, albeit at an increased cost compared to empirical potentials.
@@
Although MLPs can deliver significant accuracy improvements compared to empirical forcefields, their computational expense and the high sampling requirements of free energy calculations have meant that their application has been mostly restricted to corrective perturbations [30, 31, 32]. In this approach, the intramolecular interactions of a subset of atoms (usually the organic molecule) are modelled by the MLP, while the remaining bonded interactions and all nonbonded interactions are modelled by the empirical forcefield. Although there is some evidence that this approach can modestly improve the accuracy of RBFE calculations, it does not address the parametrisation of nonbonded interactions, which are known to be crucial for obtaining accurate free energies [31]. Indeed, it has recently been shown that, for hydration free energies, MLP corrections with mechanical embedding fail to provide statistically significant accuracy improvements [32]. It is worth noting that recently more sophisticated electrostatic embedding ML/MM schemes have also been proposed. However, the performance of these methods in corrective binding free energy calculations has not yet been established [33].
@@
In parallel, a handful of free energy methods that do not rely on alchemical transformations have been proposed, making the calculation conceptually more straightforward, easier to set up and more amenable to MLPs. One promising example uses the alchemical transfer method (ATM) [34], which relies on a coordinate transformation to interpolate between two physical end states [35, 36]. It has been shown that, especially in combination with the ANI-2x potential for modelling ligand intramolecular forces [22], this method is competitive with the commercially available FEP+ package [3], which uses the empirical OPLS4 forcefield [37]. This work still relies upon a mechanically embedded ML/MM Hamiltonian, and since both physical end states must be accommodated within the same simulation box, it is less efficient than alchemical perturbation methods due to the large amount of excess solvent required [34].
@@
Given the potential for significant accuracy improvements, there is a need for rigorous and efficient alchemical free energy methods that are compatible with (i) describing the entire system using MLPs, and (ii) existing well-tested free energy protocols and analysis packages (e.g. pymbar [38], pmx [39, 40]) that are currently in use with classical forcefields. In this work, we present a pre-trained, transferable biomolecular MLP based on the recently introduced MACE-OFF23 potentials [23] that is uniquely equipped with scalable soft-core interactions, enabling numerically stable simulations of condensed phase systems with alchemically decoupled intermolecular interactions. We provide an efficient implementation of the method in the widely used OpenMM package [8]. Furthermore, we demonstrate the fast convergence and sub-chemical accuracy of our approach in the calculation of hydration free energies for a series of small organic molecules. To the best of our knowledge, this work represents the first thermodynamically rigorous, ab initio-accurate approach to condensed phase hydration free energy calculations.
@@ II.1 Alchemical simulations with machine learned forcefields
To address the divergence of the potential energy predicted by a MLP fitted to DFT as atoms overlap when sampling intermediate values of $\lambda$, we augment the training set with dimer configurations containing artificially softened two-body interactions to avoid the formal energy divergence. We further incorporate a $\lambda$ dependence via modification of the non-trainable parameters of the model architecture to scale nonbonded interactions between the solute and the solvent. This combined approach results in $\lambda$-dependent many-body interactions that enable smooth and regularised interpolation between the coupled and decoupled end states.
@@ II.1.1 Softcore dimer curves
To construct the softcore dimer curves, we started by matching the value and gradient of the DFT force with a quartic function of the form $ax^{4}+b$ that approaches a constant value as separation goes to zero (Figure 1). The softcore energies were then obtained by integration to ensure data consistency. The transition point was adjusted for each dimer curve to minimise the gradient at the switching point, which in turn limited the maximum energy of pairwise interactions to a numerically stable value (see Figure 1). Importantly, the transition point was chosen so as not to alter the position and curvature of the energy minimum, preserving the equilibrium properties of the potential. This also ensured that the softcore region is only explored during alchemical simulations in which the nonbonded interactions are (partially) decoupled, since it lies several eV higher in energy than the equilibrium.
The sharper curvature of quartic functions enabled smaller maximum forces and energies compared to lower-order polynomials that have been employed in previous softcore formulations [16]. Using the original MACE-OFF23 dataset augmented with the synthetic dimer curve, we then fitted the softcore potential from scratch. Since it is expected that the learned radial embeddings will differ substantially from the MACE-OFF23 models due to the new softcore data, we found that this resulted in lower force and energy errors compared to transfer learning from the original MACE-OFF23 checkpoints.
Figure 1: Construction of softcore dimer curve for the Br–N diatomic pair by matching the gradient at a set switching point (in this example, 1.73 Å) to that given by DFT.
The errors of the test set on the original MACE-OFF23 data set were comparable to the performance of the original potentials, indicating that the model’s ability to learn energies and forces of equilibrium configurations is not affected by additional soft-core data (section V.1).
@@ II.1.2 Alchemical modification to MACE
Whilst in principle the softcore formulation described above is sufficient to perform stable alchemical free energy calculations, in practice, the resulting linear formulation, where the functional form of the potential has no explicit dependence on $\lambda$, can lead to significant challenges with phase space overlap [41]. Functions lacking this $\lambda$ dependence typically require complex reweighting strategies to adjust window spacing on the fly to maintain reasonable overlap between neighbouring replicas. To avoid these issues, we also modified the nonlearnable parameters in MACE to enable alchemical scaling of selected nonbonded interactions, resulting in a formulation analogous to classical soft core force fields  [15, 16, 17]. By modifying only the nonlearnable components of MACE, our approach enables alchemical simulations of arbitrary systems without additional fine tuning of the model.
Whilst we introduce alchemical modifications in terms of the MACE architecture, we stress that our approach can, in principle, be applied to any architecture that utilises the ‘density trick’ to construct atomic environment vectors without significant modification [42].
Here we describe the changes to the MACE architecture required to implement scaled nonbonded interactions. For a full description of the architecture, we direct the reader to the original publications [43, 29].
@@
In a similar approach to other recently published work [46], our formulation modifies MACE by including an additional factor of $\alpha_{ij}$ that scales selected two-body terms in the one-particle basis. These terms are selected as those that cross the alchemical boundary, for example edges connecting atoms in the solute and solvent in the case of a hydration free energy calculation.
@@
This approach (named MACE-OFF23-SC) enables smoothly scaled many-body nonbonded interactions whilst maintaining the correct physical description of the end states, where the decoupled system’s graph is identical to that of the separated components, ensuring consistent free energy calculations.
In MACE-OFF23-SC, an approximately linear $\lambda$ schedule can be applied to the alchemical transformation (Figure 2), regardless of the size of the perturbation, compared to the highly skewed schedule required for linear softcore potentials [41]. It is also worth noting that, unlike linear scaling schemes, this approach is amenable to nonequilibrium switching approaches, in which $\lambda$ is driven at a constant rate between the end states [4, 47].
@@
Figure 2: Dependence on $\lambda$ of softcore two-body interactions for a C–F diatomic pair learned by MACE-OFF23-SC.
We note that, whilst our alchemical modifications to the MACE graph share similarities with recently published work by Nam et al [46], the additional inclusion of the softcore potential is crucial for stable alchemical condensed phase simulations, where, unlike materials applications, direct atomic overlap in partially decoupled states must be handled.
II.2 Hydration free energy calculations
Figure 3: Comparison of MACE-OFF23-SC hydration free energies with classical forcefields and experiment
Hydration free energies play a crucial role in assessing the accuracy of the intermolecular interactions within a forcefield and are an essential component of the standard evaluation of thermodynamic property prediction. These simulations probe the subtle nonbonded interactions between molecules, but their sampling requirements remain small relative to, for example, protein–ligand binding free energies, hence errors can be primarily attributed to the accuracy of the forcefield.
However, unlike properties such as densities and enthalpies of vapourisation and mixing, which are directly observable as ensemble averages from MD trajectories, free energies require evaluation of expectation values such as that shown in Eq. (2). Here we employ an alternative formulation, based on Bennett’s Acceptance Ratio (BAR) [48, 38], for the free energy difference between states A and B, given by
$\Delta G_{AB}=-k_{b}T\,\ln{\left\langle e^{-\beta(U_{A}(\vec{r})-U_{B}(\vec{r}))}\right\rangle_{A}}$ (7)
[…]
These calculations can be efficiently performed using perturbations in parameter space, where the intermolecular forces are gradually decoupled between the solute and solvent. Although such free energy calculations are routine with empirical forcefields, to the best of our knowledge, this work demonstrates the first rigorous, ab initio quality hydration free energy calculations.

### v3

# Computing solvation free energies of small molecules with first principles accuracy

@@ Abstract
Free energies play a central role in characterising the behaviour of chemical systems and are among the most important quantities that can be calculated by molecular dynamics simulations. Solvation free energies in various organic solvents, in particular, are well-studied physicochemical properties of drug-like molecules and are commonly used to assess and optimise the accuracy of nonbonded parameters in empirical forcefields, and also as a fast-to-compute surrogate of performance for protein-ligand binding free energy estimation. Machine learned potentials (MLPs) show great promise as more accurate alternatives to empirical forcefields, but are not readily decomposed into physically motivated functional forms, which has thus far rendered them incompatible with standard alchemical free energy methods that manipulate individual pairwise interaction terms. However, since the accuracy of free energy calculations is highly sensitive to the forcefield, this is a key area in which MLPs have the potential to address the shortcomings of empirical forcefields. In this work, we introduce an efficient alchemical free energy protocol that enables calculations of rigorous free energy differences in condensed phase systems modelled entirely by MLPs. Using a pretrained, transferrable, alchemically equipped MLP model, we demonstrate sub-chemical accuracy for the solvation free energies of a wide range of organic molecules.
@@ I Introduction
Free energies are arguably the most important quantities accessible by molecular simulation in soft matter and play an important role in computational drug discovery [1, 2]. Free energy methods rely on rigorous thermodynamic protocols, enabling prediction of protein-ligand binding [3, 4], small molecule solvation and solubility [5], organic crystal polymorphs ranking [6] and protein residue mutation [7].
@@
Whilst various methods for calculating free energies for small molecules via MD have been proposed, alchemical transformations have emerged as the de facto standard approach, allowing free energy estimates to be obtained in a few GPU hours per compound [11, 12]. Rather than explicitly simulating a compound undergoing a change of state as it would happen in reality (which would be computationally prohibitive), alchemical calculations take advantage of the fact that the free energy is a state function and so the free energy difference between two state points is independent of the pathway connecting them. By introducing an alchemical parameter, $\lambda$, a new Hamiltonian is constructed as a linear combination of the Hamiltonians describing the two end states,
$H(\vec{r},\lambda)=\lambda H_{1}(\vec{r})+(1-\lambda)H_{0}(\vec{r})$ (1)
The free energy difference associated with the transformation can then be computed by a variety of estimators, for example by thermodynamic integration:
@@
where $\langle\dots\rangle_{\lambda}$ denotes the ensemble average corresponding to the Hamiltonian $H(\lambda)$. Using such artificial pathways leads to enormous computational savings, since the sampling required to explore intermediate state points is orders of magnitude less than it would be for a real pathway describing the physical process. In practice, the transformation is accomplished by identifying the atoms that change during the transformation (i.e. those atoms whose chemical element changes, or that disappear entirely) and interpolating the corresponding forcefield parameters between the end states. This requires the use of so-called dummy atoms to model particles that disappear during the transformation. These atoms maintain their bonded terms but have their interactions with the surroundings switched off. This allows for transformations like morphing a methyl group to a hydrogen atom or isolating a molecule from its environment, while minimally perturbing the phase space required to be explored by the system.
Early attempts to perform free energy calculations with MD faced energy conservation and convergence issues due to the divergence of the nonbonded energy (usually a combination of Lennard-Jones and Coulomb contributions) when partially decoupled atoms overlap [13, 14]. Beutler et al were the first to address this issue by incorporating softening parameters to scale nonbonded interactions as a function of the alchemical parameter $\lambda$, thus averting singularities in the potential energy as atoms come into contact [15]:
@@
where $\epsilon$ and $\sigma$ are the Lennard-Jones well depth and radius, respectively, and $\alpha_{LJ}$, $m$ and $n$ are positive, tunable constants that control the smoothness and decay of the softcore function. In the original implementation, $\alpha_{LJ}=0.5$, $n=4$ and $m=2$, although $m=n=1$ has been shown to reduce the variance of the free energy estimate [15, 16, 17]. Although Beutler-type soft core potentials are commonly used in various simulation packages, it is worth noting that several alternative approaches have been suggested to further improve numerical stability and address specific failure modes [18, 19], and alternative nonbonded functional forms with a more natural soft core have also been tried [20].
@@
While these soft-core potentials have been widely successful, a significant limitation of current state-of-the-art free energy calculations lies in the accuracy of the empirical forcefields. Specifically, the constraints imposed by the functional form, omission of energetic contributions and restricted atom-type parameters in small molecule forcefields have a significant impact on accuracy [1]. Particularly notable are the widespread use of fixed-charge electrostatic models that do not account for geometry-dependent polarization, and the limited accuracy of torsional barriers, which often require refitting to bespoke DFT calculations of each new compound to achieve reasonable accuracy [21, 22, 23].
@@
In recent years, transferable machine learned potentials (MLP)s have been introduced as a compelling alternative to empirical forcefields, demonstrating significant advances in materials modelling and biomolecular simulation [24, 25, 26, 27, 28, 29]. Building upon Behler and Parinello’s initial work [30], the field has expanded rapidly, with many architectures being proposed to accurately model the QM potential energy surface. A particularly successful innovation has been the specific encoding of permutational, rotatational and translational symmetries [31]. These approaches have led to a series of data-efficient, universal potentials for biomolecular and materials simulation, enabling accurate prediction of atomistic and thermodynamic properties for a wide range of chemical compositions, albeit at an increased cost compared to empirical potentials.
@@
Although MLPs can deliver significant accuracy improvements compared to empirical forcefields, their computational expense and the high sampling requirements of free energy calculations have meant that their application has been mostly restricted to corrective perturbations [32, 33, 34]. In this approach, the intramolecular interactions of a subset of atoms (usually the organic molecule) are modelled by the MLP, while the remaining bonded interactions and all nonbonded interactions are modelled by the empirical forcefield. Although there is some evidence that this approach can modestly improve the accuracy of RBFE calculations, it does not address the parametrisation of nonbonded interactions, which are known to be crucial for obtaining accurate free energies [33]. Indeed, it has recently been shown that, for hydration free energies, MLP corrections with mechanical embedding fail to provide statistically significant accuracy improvements [34]. It is worth noting that more sophisticated electrostatic embedding ML/MM schemes have recently been proposed [35, 36]. However, the performance of these methods in corrective binding free energy calculations has not yet been established.
@@
In parallel, a handful of free energy methods that do not rely on alchemical transformations have been proposed, making the calculation conceptually more straightforward, easier to set up and more amenable to MLPs. One promising example uses the alchemical transfer method (ATM) [37], which relies on a coordinate transformation to interpolate between two physical end states [38, 39]. It has been shown that, especially in combination with the ANI-2x potential for modelling the ligand intramolecular potential energy surface [24], this method is competitive with the commercially available FEP+ package [3], which uses the empirical OPLS4 forcefield [40]. This work still relies upon a mechanically embedded ML/MM Hamiltonian, and since both physical end states must be accommodated within the same simulation box, it is less efficient than alchemical perturbation methods due to the large amount of excess solvent required [37].
@@
Given the potential for significant accuracy improvements, there is a need for rigorous and efficient alchemical free energy methods that are compatible with (i) describing the entire system using MLPs, and (ii) existing well-tested free energy protocols and analysis packages (e.g. pymbar [41], pmx [42, 43]) that are currently in use with classical forcefields. In this work, we present a pre-trained, transferable biomolecular MLP based on the recently introduced MACE-OFF potentials [25] that is uniquely equipped with scalable soft-core interactions, enabling numerically stable simulations of condensed phase systems with alchemically decoupled intermolecular interactions. We provide an efficient implementation of the method in the widely used OpenMM package [8]. Furthermore, we demonstrate the fast convergence and sub-chemical accuracy of our approach in the calculation of solvation free energies and distribution coefficients for a series of organic molecules.
@@ II.1 Alchemical simulations with machine learned forcefields
To address the divergence of the potential energy predicted by a MLP fitted to DFT, as atoms overlap when sampling intermediate values of $\lambda$, we augmented the training set with dimer configurations containing artificially softened two-body interactions. We further incorporated a $\lambda$ dependence via modification of the non-trainable parameters of the model architecture to scale nonbonded interactions between the solute and the solvent. This combined approach results in $\lambda$-dependent many-body interactions that enable smooth and regularised interpolation between the coupled and decoupled end states. We note that, since only the nontrainable parameters are modulated during the simulation, no major changes to the training protocol are required. In all cases, the potential was trained corresponding to the $\lambda=1$ or fully interacting state, whilst the message scaling is only applied at inference time to the fitted potential.
@@ II.1.1 Softcore dimer curves
To construct the softcore dimer curves, we first computed dimer curves with DFT at the same $\omega$B97M-D3BJ/def2-TZVPPD level of theory as the SPICE dataset [44] for all combinations of the 10 elements covered by MACE-OFF [25]. For each pair, we started by matching the value and gradient of the DFT force with a polynomial of the form $ax^{10}+b$ that approaches a constant value as $r\to 0$ (Figure 1). The softcore energies were then obtained by analytical integration of the polynomial to ensure consistent energy and force data. Without appropriate precautions, this approach can lead to artificial softening of the minimum of the dimer curve if the switching point is too close to the minimum. Conversely, a switching point that is too far from equilibrium would lead to large energy and force values for small separations, which are more challenging for the model to learn.
To balance these competing factors, the transition point was adjusted for each dimer curve to minimise the gradient of the force at the switching point, down to a minimum value of 3 eV/Å2, which in turn limited the maximum energy of pairwise interactions to a numerically stable value, typically around 300 kcal/mol above the minimum. However, by setting a minimum gradient for the switching point, the transition point was chosen so as not to alter the position and curvature of the energy minimum, thereby preserving the equilibrium properties of the potential. This therefore ensures that the softcore region is only explored during alchemical simulations in which the nonbonded interactions are (partially) decoupled, since this region of the energy landscape lies several hundred kcal/mol higher in energy than equilibrium. All switching point distances are tabulated in Section S4.
The steeper curvature of tenth order functions enables smaller maximum forces and energies compared to lower-order polynomials that have been employed in previous softcore formulations [18]. Using the MACE-OFF24 training dataset [25] augmented with the synthetic dimer curves, we then fitted the softcore potential from scratch. Since it is expected that the learned radial embeddings will differ substantially from the MACE-OFF models due to the new softcore data, we found that this resulted in lower force and energy errors compared to transfer learning from the original MACE-OFF checkpoints.
Figure 1: Construction of softcore dimer curve for the Br–O diatomic pair by matching the gradient at a set switching point to that given by DFT.
The test set errors for the MACE-OFF data set were comparable to the performance of the original potentials [25], indicating that the model’s ability to learn energies and forces of equilibrium configurations is not affected by additional soft-core data (Section S1). We further confirmed that the predicted density of liquid water is unaffected by the new potential, with an average value of $1.02\pm 0.02$ g/mL at 298 K (compared to 0.98 g/mL for MACE-OFF24(M)).
It is worth noting that, whilst the introduction of softcore interactions is not expected to affect the ensemble sampled by the model at room temperature, since the curvature of the minimum is not affected and a substantial repulsive component is still learned by the model, this may not necessarily translate to high temperature regimes. In this case, the softcore region of the potential may be thermally accessible in the fully interacting state, rather than only being accessible in the partially decoupled states.
@@ II.1.2 Alchemical modification to MACE
Whilst in principle the softcore formulation described above is sufficient to perform stable alchemical free energy calculations, in practice, the resulting linear formulation, where the functional form of the potential has no explicit dependence on $\lambda$, can lead to significant challenges with phase space overlap [45]. Functions lacking this $\lambda$ dependence typically require complex reweighting strategies to adjust window spacing on the fly to maintain reasonable overlap between neighbouring replicas. To avoid these issues, we also modified the nonlearnable parameters in MACE to enable alchemical scaling of selected nonbonded interactions, resulting in a formulation analogous to classical soft-core force fields  [15, 18, 19]. By modifying only the nontrainable parameters of MACE, our approach enables alchemical simulations of arbitrary systems without additional fine tuning of the model or modification of the training protocol.
Whilst we introduce alchemical modifications in terms of the MACE architecture, we stress that our approach can be applied to any architecture that parameterises the local energy as a function of two-body terms, as has been recently demonstrated elsewhere [46]. Here we describe the changes to the MACE architecture required to implement scaled nonbonded interactions. For a full description of the architecture, we direct the reader to the original publications [47, 31].
@@
In a similar approach to other recently published work [50], our formulation modifies MACE by including an additional factor of $\alpha_{ij}$ that scales selected two-body terms in the one-particle basis. These terms are selected as those that cross the alchemical boundary, for example edges connecting atoms in the solute and solvent in the case of a solvation free energy calculation.
@@
This approach (named MACE-OFF24-SC) enables smoothly scaled many-body nonbonded interactions whilst maintaining the correct physical description of the end states, where the decoupled system’s graph is identical to that of the separated components, ensuring consistent free energy calculations.
In MACE-OFF24-SC, an approximately linear $\lambda$ schedule can be applied to the alchemical transformation (Figure 2), regardless of the size of the perturbation, compared to the highly skewed schedule required for linear softcore potentials [45]. It is also worth noting that, unlike linear scaling schemes, this approach is amenable to nonequilibrium switching approaches, in which $\lambda$ is driven at a constant rate between the end states [4, 51] (Section S5).
Figure 2: Dependence of softcore two-body interactions on $\lambda$ for a C–F diatomic pair learned by MACE-OFF24-SC.
@@
We also note that our alchemical modifications to the MACE graph share similarities with recently published work by Nam et al. [50]. In particular, this work also introduces a scaling factor on the one-particle basis functions of MACE to compute free energy changes in solids. Whilst this is sufficient for studying crystalline systems, the additional inclusion of the softcore potential is crucial for stable alchemical condensed phase simulations, where, unlike materials applications, direct atomic overlap in partially decoupled states must be handled.
II.2 Solvation free energy calculations
Solvation free energies play a crucial role in assessing the accuracy of the intermolecular interactions within a forcefield and are an essential component of the standard evaluation of thermodynamic property prediction. These simulations probe the subtle nonbonded interactions between molecules, but their sampling requirements remain small relative to, for example, protein–ligand binding free energies, hence errors can be primarily attributed to the accuracy of the forcefield.
However, unlike properties such as densities and enthalpies of vapourisation and mixing, which are directly observable as ensemble averages from MD trajectories, free energies require evaluation of expectation values such as that shown in Eq. (2). Here we employ the MBAR estimator to compute the free energy difference between the end states A and B, by combining information from a series of intermediate alchemical windows in a statistically optimal fashion [52, 41]. We stress, however, that our approach is agnostic to the estimator used.
These calculations can be efficiently performed using perturbations in parameter space, where the intermolecular interactions are gradually decoupled between the solute and solvent. Although such free energy calculations are routine with empirical forcefields, to the best of our knowledge, this work demonstrates the first rigorous, ab initio quality solvation free energy calculations with all molecules treated using MLPs.
II.2.1 Hydration free energies
We first evaluated the ability of MACE-OFF24-SC to compute accurate, converged hydration free energies for a series of organic molecules (Figure 3). For this test, we selected a diverse set of compounds from the FreeSolv database designed to cover a diverse range of functional groups relevant to medicinal chemistry [53]. Compounds were first binned according to their functional groups and subsequently filtered to remove those that were not unambiguously neutral at pH 7, as determined by Schrödinger’s Epik program [54]. This ensured compatibility with MACE-OFF24-SC, which was trained on neutral molecules only. Compounds were then randomly sampled from each bin, ensuring good functional group coverage. The identities of the selected 36 compounds are tabulated in Section S3.
We performed replica exchange molecular dynamics with an ensemble of replicas with $\lambda$ windows spaced in the interval [0,1] to connect the physical end states. To expedite phase space exploration, exchanges of replicas between adjacent $\lambda$ windows were attempted every 1 ps.
Figure 3: Comparison of MACE-OFF24-SC hydration free energies with classical forcefields and experiment. Experimental error bars are shown for those compounds where the value exceeds 0.6 kcal/mol. The shaded region represents a 1 kcal/mol deviation.
In all cases, simulations were performed with 16 replicas, which we found enabled sufficient phase-space overlap to converge the free energy simulations (Figure 5). The replica spacing was hand-tuned from an initial linear spacing to concentrate sampling in the region $0.15\leq\lambda\leq 0.4$, where the curvature of $\frac{\partial H}{\partial\lambda}$ is greatest (Section S5). For each system, convergence of the free energy as a function of time was confirmed (Figure 4 and Figure S9).